OpenStudy (anonymous):

Given sinx=-3/5 and x is in quadrant 3, what is the value of tan(x/2)?

OpenStudy (dumbcow):

sin^2 + cos^2 = 1 (-3/5)^2 + cos^2 = 1 cos^2 = 1-9/25 = 16/25 cosx = -4/5 (neg because in quad 3) Use Half angle identity for tangent tan(x/2) = sqrt((1-cos)/(1+cos)) = sqrt((1+4/5)/(1-4/5)) =sqrt(9/5)/(1/5)) = sqrt(9) = 3

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

no problem, it might be -3 though since x/2 would be in quadrant 2

OpenStudy (dumbcow):

for these problems i suggest making a trig reference sheet, thats what i do, no way i can memorize all these formulas